# RUBIES: JWST/NIRSpec Confirmation of an Infrared-luminous, Broad-line Little Red Dot with an Ionized Outflow

### v1

@@ Abstract
The JWST discovery of “little red dots” (LRDs) is reshaping our picture of the early Universe, yet the physical mechanisms driving their compact size and UV-optical colors remain elusive. Here we report an unusually bright LRD () observed as part of the RUBIES program. This LRD exhibits broad emission lines (FWHM), a blue UV continuum, a clear Balmer break and a red continuum sampled out to rest-frame 4 with MIRI. We develop a new joint galaxy and active galactic nucleus (AGN) model within the Prospector Bayesian inference framework and perform spectrophotometric modeling using NIRCam, MIRI, and NIRSpec/Prism observations. Our fiducial model reveals a galaxy alongside a dust-reddened AGN driving the optical emission. Explaining the rest-frame optical color as a reddened AGN requires , suggesting that a great majority of the accretion disk energy is re-radiated as dust emission. Yet despite clear AGN signatures, we find a surprising lack of hot torus emission, which implies that either the dust emission in this object must be cold, or the red continuum must instead be driven by a massive, evolved stellar population of the host galaxy—seemingly inconsistent with the high EW broad lines (H rest-frame EW Å). The widths and luminosities of Pa-, Pa-, Pa-, and H imply a modest black hole mass of M. Additionally, we identify a narrow blue-shifted He i m absorption feature in NIRSpec/G395M spectra, signaling an ionized outflow with kinetic energy up to % the luminosity of the AGN. The low redshift of RUBIES-BLAGN-1 combined with the depth and richness of the JWST imaging and spectroscopic observations provide a unique opportunity to build a physical model for these so-far mysterious LRDs, which may prove to be a crucial phase in the early formation of massive galaxies and their supermassive black holes.
@@ 1 Introduction
In this paper, we present a bright LRD at a spectroscopic redshift of that is luminous in the IR and has multiple broad emission lines indicative of an AGN. RUBIES-BLAGN-1 is discovered as part of the RUBIES program (JWST-GO-4233; PI de Graaff), for which we obtained high signal-to-noise spectra in both the JWST/NIRSpec low-resolution Prism and the medium-resolution G395M modes, allowing for an in-depth characterization of the physics of this object. Notably, RUBIES-BLAGN-1 is detected in JWST/MIRI F770W and F1800W imaging from PRIMER (JWST-GO-1837; PI Dunlop).
The structure of this paper is as follows. Section 2 provides an overview of the data, including imaging, selection, and spectroscopy. Section 3 details the SED modeling, emission-line decomposition, and photoionization modeling. Section 4 presents the results on the inferred AGN and host galaxy properties, gas kinematics, and mass outflow rate. We conclude in Section 5 with discussion of the key findings.
@@ 2.1 Imaging
The RUBIES targets in the UDS were selected based on public JWST/NIRCam imaging from the PRIMER survey (JWST-GO-1837; PI Dunlop), which provides NIRCam imaging in 8 bands (F090W, F115W, F150W, F200W, F277W, F356W, F444W and F410M) and MIRI imaging in the F770W and F1800W bands. Archival imaging from the Hubble Space Telescope (HST) is obtained from the CANDELS survey (Grogin et al., 2011; Koekemoer et al., 2011), which adds in 7 more bands in F435W, F606W, F814W, F105W, F125W, F140W, and F160W.
@@
We use a larger circular aperture of radius to measure fluxes from the MIRI images. The MIRI images were reduced in August 2023 (v7.0 on the DJA), prior to the release of new photometric calibrations. Following the JWST documentation, we therefore multiply the measured fluxes with a correction factor based on the updated reference file (jwst_miri_photom_0172); a factor 0.85 and 1.03 for the F770W and F1800W bands, respectively. Finally, we use the empirical PSF models of Libralato et al. (2023) to scale the MIRI aperture fluxes to total fluxes. This is motivated by our analysis that RUBIES-BLAGN-1 is unresolved in the long wavelength images. Specifically, we use pysersic (Pasha & Miller, 2023) to perform Sérsic profile fitting to the source in each long-wavelength image (F277W, F356W, F410M, F444W) shown in Fig 1, using empirical PSF models from Weibel et al. in prep. We find that the source is unresolved in all long-wavelength filters, with a half-light size smaller that a single pixel, kpc. For the remainder of the paper, we conservatively use an upper limit for the half-light radius of 1 pixel (0.3 kpc) for all comparisons involving the half-light radius.
@@ 3.1.3 AGN Accretion Disk
We take the combination of the above AGN accretion disk and torus model and the galaxy model (§ 3.1.2) as the fiducial model of this paper. The relevant parameters are listed in Table 1. However, the unknown nature of LRDs means that other possibilities cannot be ruled out a-priori. We thus additionally fit the data with three alternative models, including fitting with AGN light only, without the new dust component, and fitting the data without the UV continuum. We use these other models to infer the systematic errors in our inferred parameters, and to propose discerning hypotheses for future testing. Details are presented in Appendix A.
@@
ratio between the AGN and the host galaxy’s flux densities at rest-frame 1500 Å, pre-attentuation |
log uniform: , |
ratio between the AGN luminosity in the mid-IR and the host galaxy’s bolometric luminosity |
log uniform: , |
optical depth of AGN torus dust |
log uniform: , |
power law index for a Calzetti et al. (2000) attenuation curve |
uniform: , |
optical depth of diffuse dust seen by both the host galaxy and the AGN |
uniform: , |
power law index for the AGN-only attenuation curve |
uniform: , |
optical depth of dust seen only by the AGN |
uniform: , |
fraction of starlight that is not attenuated by the diffuse dust |
uniform: , |
@@ 3.1.5 Emission line Marginalization
3.2 Emission line Decomposition
@@
We leverage the data from the two independent visits for a better sampling of the LSF. Instead of combining the two spectra, which may introduce further correlated noise, we simultaneously fit to the two spectra. We use the Markov Chain Monte Carlo ensemble sampling method implemented in emcee (Foreman-Mackey et al., 2013) to estimate the posterior distributions of the model parameters.
@@
To fit the H line we additionally include the [N ii] doublet in the model. We assume that the [N ii] lines are narrow and equal to the narrow line width of H, and fix the flux ratio of the doublet to 1:2.94. Because the narrow line width is a factor narrower than the LSF, we cannot constrain the width from the prism data alone. We therefore assume a Gaussian prior for the velocity dispersion of the narrow line based on the estimates from the well-resolved Paschen lines: and .
@@ 3.4 Photoionization Modeling
We estimate the hydrogen column density associated with the He i absorber by running photoionization models using the C23 version of cloudy (Chatzikos et al., 2023). We adopt the standard AGN radiation field which is a modified version of the Mathews & Ferland (1987) SED with a sub-millimeter break at 10 microns. We calculate a two-dimensional grid of models with hydrogen number density varied in the range 2 [cm] 10 in 0.5 dex steps and ionization parameter varied in the range 0 in 0.3 dex steps. The elemental abundances are fixed to solar. The models are run until they reach a hydrogen column density of [cm] = 24.
@@ 4 Results
With those ideas in mind, we begin by presenting results from fitting only starlight to the rest UV/optical continuum of RUBIES-BLAGN-1 (§4.1). This model serves as a useful benchmark, against which we evaluate our AGN and galaxy composite model (§4.2). Finally, we explore the family of models that are consistent with the mid-to-far infrared data and limits, particularly in the context of dusty torus models that characterize typical AGNs (§4.3). Parameter constraints, including inferred properties of the host galaxy and the AGN are listed in Table 2, and line fluxes and kinematics are listed in Table 3.
@@
, H | 8.6 (8.5)ddThe black hole (BH) masses in the first column are estimated assuming the fiducial model, and those in parentheses are based on an alternative dust model (see Appendix A). |
, Pa | 8.4 (8.3)ddThe black hole (BH) masses in the first column are estimated assuming the fiducial model, and those in parentheses are based on an alternative dust model (see Appendix A). |
, Pa | 7.9 (7.8)ddThe black hole (BH) masses in the first column are estimated assuming the fiducial model, and those in parentheses are based on an alternative dust model (see Appendix A). |
, Pa | 7.9 (7.8)ddThe black hole (BH) masses in the first column are estimated assuming the fiducial model, and those in parentheses are based on an alternative dust model (see Appendix A). |
@@ 4.1 Galaxy-only Best Fit
As seen from Figure 3, the galaxy-only model produces a good fit in the rest-frame UV-optical and an inferred star mass of with an sSFR of . The rest of the inferred properties are listed in Appendix A, including its mass-weighted age of Gyr and of .
It would be unusual but not impossible to have such an old massive galaxy (e.g., the quiescent galaxy at , which has in the center; Setton et al. 2024). It would also be surprising to have such strong broad-line features (H rest-frame EW Å) and such a massive black hole with very little continuum contribution from the accretion disk, although AGN broad lines may still exist in older galaxies (Carnall et al., 2023). Moreover, the small size of RUBIES-BLAGN-1 ( kpc, conservatively using an upper limit to half-light radius of pixel; see Section 2) implies a very high stellar density (Baggen et al., 2023). This size would alao make RUBIES-BLAGN-1 a small galaxy relative to its peers, but not an outlier among known old galaxies at similar mass and redshift (Wright et al., 2023; Ji et al., 2024). However, the inferred narrow line widths of the Pa- and Pa- lines () and small size imply a dynamical mass (where is the gravitational constant; see e.g., Cappellari et al., 2006), which would be difficult to reconcile with a stellar mass estimate of even when considering projection effects that may underestimate the true velocity dispersion of the system. We therefore favor the fiducial galaxy and AGN joint fit presented in the following section, but cannot definitively rule out a galaxy-dominated continuum. Follow-up deep medium-resolution data to reveal possible stellar absorption lines, or deeper rest-frame m data, will help to discriminate the different scenarios.
@@ 4.2 UV/Optical Best Fit
In Figure 4 we show our fiducial galaxy and AGN joint model. An immediate observation is the capability of our fiducial model to generate a very red slope in the rest optical region and a blue slope in the rest near-UV, providing an effective fit to the observed data. This v-shape in is difficult to reproduce with a simple AGN-only model (see Appendix A). As depicted in Figure 4, starlight makes up the rest-UV, whereas the light from the reddened AGN accretion disk dominates the spectrum red-ward of rest 4000 Å.
@@
The preference for a metal-poor solution is primarily caused by the need to predict the UV excess, since the fiducial model essentially couples the UV light to the galaxy. A natural question is thus whether the host galaxy properties are heavily influenced by the UV continuum. We address this question by fitting our fiducial model to the rest optical part of the spectrum only in Appendix A. In brief, we find that the inferred stellar mass as well as the inferred star-formation history in this case are similar to those from the fiducial model, suggesting that the UV continuum is not driving the inferred star formation history or mass.
@@ 4.4 Gas Kinematics
We show the decomposition into narrow and broad components (§ 3.2) in the Paschen and H lines in Figure 6, and list the fluxes and velocity dispersions in Table 3. We find that the narrow and broad line widths of the Pa- and Pa- lines are in good agreement, and that the narrow line flux comprises 14% and 10% of the total line flux, respectively. The narrow and broad line widths of the Pa- emission are broader than the other Paschen lines by approximately , and with a higher narrow-to-total flux ratio (25%). Although some physical differences may be expected if the different line transitions trace gas in different parts of the broad line region (BLR), the Pa- line is at the very edge of the spectrum which may bias the kinematic fitting. The H broad line is significantly broader, by a factor , than the Paschen broad lines; the narrow line flux is approximately 30% of the total H line flux. This may imply that the H line traces a closer-in region of the BLR compared to the Paschen lines (see e.g. Kim et al., 2010). However, as the narrow H line and the [N ii] lines are unresolved at the resolution of the Prism, the kinematic fit depends on the model assumptions made (at the level for the broad line width). Follow-up spectroscopy at higher resolution will be crucial to perform more careful modeling of the H and [N ii] emission line complex and robustly constrain the narrow-to-total line flux ratio, and to confirm the difference in the broad line width between the Balmer and Paschen lines.
Interestingly, we find a blue-shifted absorption feature in the wings of the He i 1.083 m line. The He i absorption of RUBIES-BLAGN-1 shows up in both G395M spectra, ruling out the possibility of a detector defect. We measure a velocity dispersion of , and a velocity offset of with respect to the narrow HeI emission. The He i line is associated with the 2S metastable state, which is 19.7 eV above the ground state and is primarily populated by recombination from He ii. Therefore, the absorption line effectively traces ionized gas. This makes it likely that the absorption tracers an outflow from RUBIES-BLAGN-1 rather than intervening gas along the line of sight. In addition, the velocity of the absorber also suggests it is associated with the outflow (e.g., Culliton et al. 2019).

### v2

@@ Abstract
The JWST discovery of “little red dots” (LRDs) is reshaping our picture of the early Universe, yet the physical mechanisms driving their compact size and UV-optical colors remain elusive. Here we report an unusually bright LRD () observed as part of the RUBIES program. This LRD exhibits broad emission lines (FWHM), a blue UV continuum, a clear Balmer break and a red continuum sampled out to rest-frame 4 with MIRI. We develop a new joint galaxy and active galactic nucleus (AGN) model within the Prospector Bayesian inference framework and perform spectrophotometric modeling using NIRCam, MIRI, and NIRSpec/Prism observations. Our fiducial model reveals a galaxy alongside a dust-reddened AGN driving the optical emission. Explaining the rest-frame optical color as a reddened AGN requires , suggesting that a great majority of the accretion disk energy is re-radiated as dust emission. Yet, despite clear AGN signatures, we find a surprising lack of hot torus emission, which implies that either the dust emission in this object must be cold, or the red continuum must instead be driven by a massive, evolved stellar population of the host galaxy—seemingly inconsistent with the high EW broad lines (H rest-frame EW Å). The widths and luminosities of Pa-, Pa-, Pa-, and H imply a modest black hole mass of M⊙. Additionally, we identify a narrow blue-shifted He i m absorption feature in NIRSpec/G395M spectra, signaling an ionized outflow with kinetic energy up to % the luminosity of the AGN. The low redshift of RUBIES-BLAGN-1 combined with the depth and richness of the JWST imaging and spectroscopic observations provide a unique opportunity to build a physical model for these so-far mysterious LRDs, which may prove to be a crucial phase in the early formation of massive galaxies and their supermassive black holes.
@@ 1 Introduction
In this paper, we present a bright LRD at a spectroscopic redshift of that is luminous in the IR and has multiple broad emission lines indicative of an AGN. RUBIES-BLAGN-1 is discovered as part of the RUBIES program (JWST-GO-4233; PIs de Graaff & Brammer), for which we obtained high signal-to-noise spectra in both the JWST/NIRSpec low-resolution Prism and the medium-resolution G395M modes, allowing for an in-depth characterization of the physics of this object. Notably, RUBIES-BLAGN-1 is detected in JWST/MIRI F770W and F1800W imaging from PRIMER (JWST-GO-1837; PI Dunlop).
The structure of this paper is as follows. Section 2 provides an overview of the data, including imaging, selection, and spectroscopy. Section 3 details the SED modeling, emission-line decomposition, and photoionization modeling. Section 4 presents the results on the inferred AGN and host galaxy properties, morphology, gas kinematics, and mass outflow rate. We conclude in Section 5 with discussion of the key findings.
@@ 2.1 Imaging
The RUBIES targets in the UDS were selected based on public JWST/NIRCam imaging from the PRIMER survey, which provides NIRCam imaging in 8 bands (F090W, F115W, F150W, F200W, F277W, F356W, F444W and F410M) and MIRI imaging in the F770W and F1800W bands. Archival imaging from the Hubble Space Telescope (HST) is obtained from the CANDELS survey (Grogin et al., 2011; Koekemoer et al., 2011), which adds in 7 more bands in F435W, F606W, F814W, F105W, F125W, F140W, and F160W.
@@
We use a larger circular aperture of radius to measure fluxes from the MIRI images. The MIRI images were reduced in August 2023 (v7.0 on the DJA), prior to the release of new photometric calibrations. Following the JWST documentation, we therefore multiply the measured fluxes with a correction factor based on the updated reference file (jwst_miri_photom_0172); a factor 0.85 and 1.03 for the F770W and F1800W bands, respectively. Finally, we use the empirical PSF models of Libralato et al. (2023) to scale the MIRI aperture fluxes to total fluxes. This is motivated by our morphological analysis (detailed in Appendix A) which shows that RUBIES-BLAGN-1 is unresolved in the long wavelength images, with a half-light size smaller that a single pixel, kpc.
@@ 3.1.3 AGN Accretion Disk
We take the combination of the above AGN accretion disk and torus model and the galaxy model (§ 3.1.2) as the fiducial model of this paper. The relevant parameters are listed in Table 1. However, the unknown nature of LRDs means that other possibilities cannot be ruled out a-priori. We thus additionally fit the data with three alternative models, including fitting with AGN light only, without the new dust component, and fitting the data without the UV continuum. We use these other models to infer the systematic errors in our inferred parameters, and to propose discerning hypotheses for future testing. Details are presented in Appendix B.
@@
ratio between the AGN and the host galaxy’s flux densities at rest-frame 1500 Å, pre-attentuation | log uniform: , | |
ratio between the AGN luminosity in the mid-IR and the host galaxy’s bolometric luminosity | log uniform: , | |
optical depth of AGN torus dust | log uniform: , | |
power law index for a Calzetti et al. (2000) attenuation curve | uniform: , | |
optical depth of diffuse dust seen by both the host galaxy and the AGN | uniform: , | |
power law index for the AGN-only attenuation curve | uniform: , | |
optical depth of dust seen only by the AGN | uniform: , | |
fraction of starlight that is not attenuated by the diffuse dust | uniform: , |
@@ 3.1.5 Emission line Marginalization
3.2 Emission-line Decomposition
@@
We leverage the data from the two independent visits for a better sampling of the LSF. Instead of combining the two spectra, which may introduce further correlated noise, we simultaneously fit to the two spectra. We use the Markov Chain Monte Carlo (MCMC) ensemble sampling method implemented in emcee (Foreman-Mackey et al., 2013) to estimate the posterior distributions of the model parameters.
@@
To fit the H line we additionally include the [N ii]6549,6585 doublet in the model. We assume that the [N ii] lines are narrow and equal to the narrow line width of H, and fix the flux ratio of the doublet to 1:2.94. Because the narrow line width is a factor narrower than the LSF, we cannot constrain the width from the prism data alone. We therefore assume a Gaussian prior for the velocity dispersion of the narrow line based on the estimates from the well-resolved Paschen lines: and .
@@ 3.4 Photoionization Modeling
We estimate the hydrogen column density associated with the He i absorber by running photoionization models using the C23 version of Cloudy (Chatzikos et al., 2023). We adopt the standard AGN radiation field which is a modified version of the Mathews & Ferland (1987) SED with a sub-millimeter break at 10 microns. We calculate a two-dimensional grid of models with hydrogen number density varied in the range 2 [cm-3] 10 in 0.5 dex steps and ionization parameter varied in the range 0 in 0.3 dex steps. The elemental abundances are fixed to solar. The models are run until they reach a hydrogen column density of [cm-2] = 24.
@@ 4 Results
With those ideas in mind, we begin by reporting the size measurements in the long and short wavelengths (§4.1). These may shed light on the AGN or galaxy origins as RUBIES-BLAGN-1 is better resolved owing to its lower redshift. We then proceed to present results from fitting only starlight to the rest UV/optical continuum of RUBIES-BLAGN-1 (§4.2). This model serves as a useful benchmark, against which we evaluate our AGN and galaxy composite model (§4.3). Finally, we explore the family of models that are consistent with the mid-to-far IR data and limits, particularly in the context of dusty torus models that characterize typical AGNs (§4.4). Parameter constraints, including inferred properties of the host galaxy and the AGN are listed in Table 2, and line fluxes and kinematics are listed in Table 3.
4.1 Sizes in the UV and Optical
RUBIES-BLAGN-1 is unresolved in the long wavelength filters, corresponding to the rest-frame optical emission, with an inferred effective radius of . However, comparison of the 1D radial profile to the PSF profile as well as 2D Sérsic profile fitting in the F115W filter, suggest that the source is marginally resolved in the rest-frame UV with . Details on the morphological modeling can be found in Appendix A.
Critically, this analysis is independent of the SED modeling, and already suggests that the UV and rest-frame optical trace different physical components.
@@ 4.1 Sizes in the UV and Optical
, H | 8.6 (8.5)ddThe black hole (BH) masses in the first column are estimated assuming the fiducial model, and those in parentheses are based on an alternative dust model (see Appendix B). |
, Pa | 8.4 (8.3)ddThe black hole (BH) masses in the first column are estimated assuming the fiducial model, and those in parentheses are based on an alternative dust model (see Appendix B). |
, Pa | 7.9 (7.8)ddThe black hole (BH) masses in the first column are estimated assuming the fiducial model, and those in parentheses are based on an alternative dust model (see Appendix B). |
, Pa | 7.9 (7.8)ddThe black hole (BH) masses in the first column are estimated assuming the fiducial model, and those in parentheses are based on an alternative dust model (see Appendix B). |
@@ 4.2 Galaxy-only Best Fit
As seen from Figure 3, the galaxy-only model produces a good fit in the rest-frame UV-optical and an inferred star mass of with an sSFR of . The rest of the inferred properties are listed in Appendix B, including its mass-weighted age of Gyr and of .
It would be unusual but not impossible to have such an old massive galaxy (e.g., the quiescent galaxy at , which has in the center; Setton et al. 2024). It would also be surprising to have such strong broad-line features (H rest-frame EW Å) and such a massive black hole with very little continuum contribution from the accretion disk, although AGN broad lines may still exist in older galaxies (Carnall et al., 2023). Moreover, the small size of RUBIES-BLAGN-1 (effective radius of kpc in the UV, or kpc conservatively from an upper limit of 1 pixel) implies a very high stellar density (Baggen et al., 2023). This size would also make RUBIES-BLAGN-1 a small galaxy relative to its peers, but not an outlier among known old galaxies at similar mass and redshift (Wright et al., 2023; Ji et al., 2024). However, the inferred narrow line widths of the Pa- and Pa- lines () and small size imply a dynamical mass (where is the gravitational constant; see e.g., Cappellari et al., 2006), which would be difficult to reconcile with a stellar mass estimate of even when considering projection effects that may underestimate the true velocity dispersion of the system.
Therefore, we favor the fiducial galaxy and AGN joint fit presented in the following section, but cannot definitively rule out a galaxy-dominated continuum. Follow-up deep medium-resolution data to reveal possible stellar absorption lines, or deeper rest-frame m data, will help to discriminate the different scenarios.
@@ 4.3 UV/Optical Best Fit
In Figure 4 we show our fiducial galaxy and AGN joint model. An immediate observation is the capability of our fiducial model to generate a very red slope in the rest optical region and a blue slope in the rest near-UV, providing an effective fit to the observed data. This v-shape in is difficult to reproduce with a simple AGN-only model (see Appendix B). As depicted in Figure 4, starlight makes up the rest-UV, whereas the light from the reddened AGN accretion disk dominates the spectrum red-ward of rest 4000 Å.
@@
The preference for a metal-poor solution is primarily caused by the need to predict the UV excess, since the fiducial model essentially couples the UV light to the galaxy. A natural question is thus whether the host galaxy properties are heavily influenced by the UV continuum. We address this question by fitting our fiducial model to the rest optical part of the spectrum only in Appendix B. In brief, we find that the inferred stellar mass as well as the inferred star-formation history in this case are similar to those from the fiducial model, suggesting that the UV continuum is not driving the inferred star formation history or mass.
Notably, the above results from SED fitting are consistent with the size measurements: the emission in the rest-frame UV originates from the host galaxy and thus is extended spatially, whereas the rest-frame optical emission is dominated by the AGN that is point-like in nature.
@@ 4.5 Gas Kinematics
We show the decomposition into narrow and broad components (§ 3.2) in the Paschen and H lines in Figure 6, and list the fluxes and velocity dispersions in Table 3. We find that the narrow and broad line widths of the Pa- and Pa- lines are in good agreement, and that the narrow line flux comprises 14% and 10% of the total line flux, respectively. The narrow and broad line widths of the Pa- emission are broader than the other Paschen lines by approximately , and with a higher narrow-to-total flux ratio (25%). Although some physical differences may be expected if the different line transitions trace gas in different parts of the broad line region (BLR), the Pa- line is at the very edge of the spectrum which may bias the kinematic fitting. The H broad line is significantly broader, by a factor , than the Paschen broad lines; the narrow line flux is approximately 30% of the total H line flux. This may imply that the H line traces a closer-in region of the BLR compared to the Paschen lines (see e.g., Kim et al., 2010). However, as the narrow H line and the [N ii] lines are unresolved at the resolution of the Prism, the kinematic fit depends on the model assumptions made (at the level for the broad line width). Follow-up spectroscopy at higher resolution will be crucial to perform more careful modeling of the H and [N ii] emission line complex and robustly constrain the narrow-to-total line flux ratio, and to confirm the difference in the broad line width between the Balmer and Paschen lines.
Interestingly, we find a blue-shifted absorption feature in the wings of the He i 1.083 m line. The He i absorption of RUBIES-BLAGN-1 shows up in both G395M spectra, ruling out the possibility of a detector defect. We measure a velocity dispersion of , and a velocity offset of with respect to the narrow HeI emission. The He i line is associated with the 23S metastable state, which is 19.7 eV above the ground state and is primarily populated by recombination from He ii. Therefore, the absorption line effectively traces ionized gas. This makes it likely that the absorption tracers an outflow from RUBIES-BLAGN-1 rather than intervening gas along the line of sight. In addition, the velocity of the absorber also suggests it is associated with the outflow (e.g., Culliton et al. 2019).